# Two-dimensional RMSD projections for reaction path visualization and validation

### 17 Feb 2026 v3

@@ Abstract
Transition state or minimum energy path finding methods constitute a routine component of the computational chemistry toolkit. Standard analysis involves trajectories conventionally plotted in terms of the relative energy to the initial state against a cumulative displacement variable, or the image number. These dimensional reductions obscure structural rearrangements in high dimensions and are often history dependent. This precludes the ability to compare optimization histories of different methods beyond the number of calculations, time taken, and final saddle geometry. We present a method mapping trajectories onto a two-dimensional projection defined by a permutation corrected root mean square deviation from the reactant and product configurations. Energy is represented as an interpolated color-mapped surface constructed from all optimization steps using a gradient aware derivative Gaussian Process. This representation highlights optimization trajectories, identifies endpoint basins, and diagnoses convergence concerns invisible in one-dimensional profiles. We demonstrate the framework on a cycloaddition reaction, showing that a machine-learned potential saddle and density functional theory reference lie on comparable energy contours despite geometric displacements, along with the ratification of the visualization for more complex reactions, a grignard rearrangement, and a bicyclobutadiene rearrangement.
@@
The first order saddle point, is often the first approximation to understand the reaction kinetics of any system of interest. Starting from two known configurations, a series of configurations connected through fictitious springs, that is, the nudged elastic band method [18] is among the most common. The NEB may be coupled with the climbing image [16] and spring variations [1] with machine learned local accelerations [14, 20, 11] and the string family of methods [28, 34, 24]. Together, these form the current Pareto optimal forms to find a saddle.
Beyond difficulties of convergence and the intricacies of the initial points chosen, the resulting path histories are best studied through the “eye-ball norm”, wherein manual inspection of the final path, along with a normal mode analysis for the saddle point estimate are considered sufficient. In turn, this means the most common visual plots, are one dimensional “profile” plots. These may show only the final optimized path, against the image number 111e.g. as found in ASE [21], which obscures any distance measure between the images. Alternatively, a well-known improvement to this involves the cubic Hermite interpolation [16] involving the forces relative to the “reaction coordinate,” defined by the piece-wise sum of the Euclidiean distances between intermediate images. The reaction coordinate $s_{i}=\sum_{j=1}^{i}\|\mathbf{R}_{j}-\mathbf{R}_{j-1}\|_{2}$ depends entirely on the specific path geometry and optimization history. This scalar measure lacks a unique definition and collapses global geometric information onto a single, arbitrary axis. This dimensional reduction precludes rigorous comparison between trajectories generated by differing algorithms, such as NEB versus Frozen String methods, or even identical algorithms with varying parameters. Furthermore, the one-dimensional projection frequently masks the distinction between numerical instability and physical relaxation into alternative basins, creating ambiguity in the validation of stationary points which costs researchers often days of calculations to clarify.
@@ 1 Methodology
Figure 1: Overview of the 2D RMSD projection pipeline. NEB output in $\mathbb{R}^{3N}$ is projected to intrinsic $(r,p)$ coordinates via IRA-based permutation-invariant RMSD. Tangential forces are projected onto the 2D tangent to construct synthetic gradients. Energies and gradients jointly feed a gradient-enhanced GP with the IMQ kernel to produce the interpolated energy surface $E(r,p)$.
The complete methodology is summarized in Figure 1. We describe the one-dimensional profiles first. Since standard NEB implementations provide discrete images $\mathbf{X}_{i}$ and energies $E_{i}$. To reconstruct a physically consistent energy profile $E(s)$ along the path, we utilize the available force information. We define the discrete reaction coordinate $s_{i}$ as the cumulative Euclidean distance:
@@ 1.1 Intrinsic Projection Coordinates
To map the high-dimensional optimization trajectory onto a consistent 2D subspace, we define the projection coordinates $(u,v)$ as the distances from the reactant ($\mathbf{R}$) and product ($\mathbf{P}$) reference configurations. To ensure the metric remains robust across automated workflows where atom indexing may vary, we employ a permutation-invariant Root Mean Square Deviation (RMSD) with optimal permutation $\mathbf{P}$ and rotation $\mathbf{R}$ determined via the Iterative Rotations and Assignments (IRA) [15] algorithm. This procedure guarantees unique, invariant coordinates regardless of the initial atom indexing or frame orientation [11]. The discrete optimization steps provide a sparse sampling of this geometric subspace.
@@
Here, $\|\cdot\|_{F}$ denotes the Frobenius norm. $\mathbf{Q}\in SO(3)$ represents the optimal rotation matrix, and $\mathbf{\Pi}$ represents the optimal permutation matrix. We solve for $\mathbf{Q}$ and $\mathbf{\Pi}$ simultaneously. This ensures that the resulting coordinates $(u,v)=(d(\mathbf{X},\mathbf{R}),d(\mathbf{X},\mathbf{P}))$ remain invariant to rigid body rotation, translation, and arbitrary atom index labeling.
[…]
@@
For systems with well-established collective variables (Ramachandran angles for peptides, donor-acceptor distances for proton transfer), such coordinates naturally provide superior physical insight [22]. However, the vast majority of transition state searches lack such convenient coordinates. Surface reactions, where an adatom hops into a subsurface site on a metallic slab, or catalytic processes involving concerted rearrangements of adsorbates, have no obvious low-dimensional collective variable. Constructing a two-dimensional free energy surface for such systems requires extensive sampling and a priori knowledge of the relevant degrees of freedom. In contrast, the RMSD-based projection requires only the Cartesian coordinates already produced by the calculation. Unlike free energy surface methods that mandate pre-specification of reaction coordinates, our approach visualizes arbitrary path based methods post hoc, requiring only the endpoints. This eliminates the circular dependency where one must already understand the mechanism to choose coordinates capable of revealing it.
@@
Our method targets the complementary regime to global PES exploration techniques: we provide geometric validation for individual optimization trajectories ($\approx 10^{3}$ samples per NEB calculation) rather than exhaustive basin catalogs requiring $\approx 10^{6}-10^{9}$ samples. Most single NEB calculations involve only $\approx 10^{3}$ geometries concentrated along a one-dimensional path connecting a single reactant-product pair. This sampling is insufficient and inappropriate for techniques designed to map complete basin connectivity or construct disconnectivity graphs [33]. Our framework instead answers a distinct question: “Did this specific optimization find a physically reasonable barrier topology?” This diagnostic capability proves critical for MLIP validation, where the goal is verifying that a machine-learned potential reproduces the correct transition pathway, not exhaustively cataloging all possible pathways.
@@ 1.2 Energy Landscape Projection
Figure 2: Dimensional reduction from the full 3N13N{-}1 configuration space to 1D (standard reaction coordinate $s$, left) versus 2D RMSD projection (this work, right). Both maps are many-to-one: distinct Cartesian configurations can share the same projected coordinates. The 1D projection loses all geometric context. The 2D projection retains endpoint distances, enabling cross-method comparison and landscape visualization, at the cost of 3N33N{-}3 unconstrained degrees of freedom.
The discrete optimization steps provide a sparse, unstructured sampling of the 2D $(u,v)$ domain (Figure 2). To visualize slices of the underlying potential energy landscape $E(u,v)$, in a meaningful way to compare data from disparate calculations, we construct a continuous projection approximating the potential energy surface. Both the 1D and 2D projections are lossy: multiple distinct configurations in $\mathbb{R}^{3N}$ can map to the same projected coordinate. The 2D projection retains strictly more information than the 1D reaction coordinate, but it is not and cannot be the full PES.
@@
Only the tangential force component $F_{\parallel}$ enters this construction. The full Cartesian force $\mathbf{F}_{i}\in\mathbb{R}^{3N}$ contains orthogonal components encoding local curvature of the true PES. However, projecting these into the $(r,p)$ subspace requires the Jacobian $\partial d_{\text{RMSD}}/\partial\mathbf{R}$, which is not analytically available: the RMSD involves a joint minimization over rotations $\mathbf{Q}\in SO(3)$ and permutations $\mathbf{\Pi}$, making the map from $\mathbb{R}^{3N}$ to $\mathbb{R}^{2}$ piecewise smooth at best. The tangential projection is therefore the consistent choice without additional approximations. It still carries non-trivial 2D information because the tangent direction $(\tau_{r},\tau_{p})$ varies along the path, distributing force information across both coordinates in a way not accessible from scalar reaction coordinate plots.
We stress that the interpolated surface is a projected slice of the full \$3N\$-dimensional PES: many distinct Cartesian configurations map to the same $(r,p)$ point. The interpolation represents the conditional expectation of the energy given the RMSD coordinates, marginalizing over the unconstrained 3N23N-2 degrees of freedom. Features of the projected surface, such as apparent alternative pathways in the color map, do not necessarily correspond to physical pathways in the full configuration space. Similarly, the property that MEP images are minima in the orthogonal hyperplane holds in the full-dimensional PES but has no analogue in a two-dimensional lossy projection. The interpolation is reliable near the sampled data and degrades with distance, as with any regression model.
@@
The hyperparameter optimization follows the subset optimization [12] concept from the OT-GPD, here, the points from the final path are used to calculate the length and noise scales, which are subsequently applied while fitting the entire surface. This subsampled approach reduces computational cost from $O(N^{3})$ to $O(n^{3})$ where $n\ll N$ (typically $n\approx 20$ vs $N\approx 500-2000$), while simultaneously improving robustness by isolating hyperparameter learning from transient optimization dynamics.
The resulting interpolated surface enables direct overlay of reference structures (e.g., DFT-optimized saddle points) onto MLIP-generated landscapes. Visual assessment of whether a potential captures qualitatively correct barrier topology becomes possible even when geometric displacements occur, a capability impossible with 1D energy profiles where the reaction coordinate axis itself depends on path geometry. This addresses a critical gap in current MLIP validation workflows.

### 24 Feb 2026 v4

@@ Abstract
Transition state or minimum energy path finding methods constitute a routine component of the computational chemistry toolkit. Standard analysis involves trajectories conventionally plotted in terms of the relative energy to the initial state against a cumulative displacement variable, or the image number. These dimensional reductions obscure structural rearrangements in high dimensions and are often history dependent. This precludes the ability to compare optimization histories of different methods beyond the number of calculations, time taken, and final saddle geometry. We present a method mapping trajectories onto a two-dimensional projection defined by a permutation corrected root mean square deviation from the reactant and product configurations. Energy is represented as an interpolated color-mapped surface constructed from all optimization steps using a gradient-enhanced Gaussian Process with the inverse multiquadric kernel, whose posterior variance contours delineate data-supported regions from extrapolated ones. A rotated coordinate frame decomposes the RMSD plane into reaction progress and orthogonal distance. We show the utility of the framework on a cycloaddition reaction, where a machine-learned potential saddle and density functional theory reference lie on comparable energy contours despite geometric displacements, along with the ratification of the visualization for more complex reactions, a Grignard rearrangement, and a conrotatory bicyclobutane ring opening.
@@
The first order saddle point is often the first approximation to understand the reaction kinetics of any system of interest. Starting from two known configurations, a series of configurations connected through fictitious springs, that is, the nudged elastic band method [19] is among the most common. The NEB may be coupled with the climbing image [17] and spring variations [1] with machine learned local accelerations [15, 21, 12] and the string family of methods [29, 36, 25]. Together, these constitute the current Pareto front for saddle searches.
Beyond convergence difficulties and the sensitivity to initial points, the resulting path histories are best studied through the “eye-ball norm”, wherein manual inspection of the final path, along with a normal mode analysis for the saddle point estimate are considered sufficient. In turn, this means the most common visual plots are one dimensional “profile” plots. These may show only the final optimized path, against the image number 111e.g. as found in ASE [22], which obscures any distance measure between the images. Alternatively, a well-known improvement to this involves the cubic Hermite interpolation [17] involving the forces relative to the “reaction coordinate,” defined by the piece-wise sum of the Euclidean distances between intermediate images. The reaction coordinate $s_{i}=\sum_{j=1}^{i}\|\mathbf{R}_{j}-\mathbf{R}_{j-1}\|_{2}$ depends entirely on the specific path geometry and optimization history. This scalar measure lacks a unique definition and collapses global geometric information onto a single, arbitrary axis. This dimensional reduction precludes rigorous comparison between trajectories generated by differing algorithms, such as NEB versus Frozen String methods [25], or even identical algorithms with varying parameters. The one-dimensional projection frequently masks the distinction between numerical instability and physical relaxation into alternative basins. The resulting ambiguity in stationary point validation costs researchers days of calculations to clarify.
@@ 1 Methodology
[…]
Figure 1: Overview of the 2D RMSD projection pipeline. NEB output in $\mathbb{R}^{3N}$ is projected to intrinsic $(r,p)$ coordinates via IRA-based permutation-invariant RMSD. Tangential forces are projected onto the 2D tangent to construct synthetic gradients. Energies and gradients jointly feed a gradient-enhanced GP with the IMQ kernel to produce the interpolated energy surface $E(r,p)$. A final rigid rotation aligns the $(r,p)$ frame to reaction progress $s$ and orthogonal deviation $d$, yielding the $E(s,d)$ landscape.
The complete methodology is summarized in Figure 1. We describe the one-dimensional profiles first. Standard NEB implementations provide discrete images $\mathbf{X}_{i}$ and energies $E_{i}$. To reconstruct a physically consistent energy profile $E(s)$ along the path, we use the available force information. We define the discrete reaction coordinate $s_{i}$ as the cumulative Euclidean distance:
@@ 1.1 Intrinsic Projection Coordinates
To map the high-dimensional optimization trajectory onto a consistent 2D subspace, we define the projection coordinates $(r,p)$ as the distances from the reactant ($\mathbf{R}$) and product ($\mathbf{P}$) reference configurations. To ensure the metric remains robust across automated workflows where atom indexing may vary, we employ a permutation-invariant Root Mean Square Deviation (RMSD) with optimal permutation $\boldsymbol{\Pi}$ and rotation $\mathbf{Q}$ determined via the Iterative Rotations and Assignments (IRA) [16] algorithm. This procedure guarantees unique, invariant coordinates regardless of the initial atom indexing or frame orientation [12]. The discrete optimization steps provide a sparse sampling of this geometric subspace.
@@
Here, $\|\cdot\|_{F}$ denotes the Frobenius norm. $\mathbf{Q}\in SO(3)$ represents the optimal rotation matrix, and $\mathbf{\Pi}$ represents the optimal permutation matrix. We solve for $\mathbf{Q}$ and $\mathbf{\Pi}$ simultaneously. This ensures that the resulting coordinates $(r,p)=(d(\mathbf{X},\mathbf{R}),d(\mathbf{X},\mathbf{P}))$ remain invariant to rigid body rotation, translation, and arbitrary atom index labeling.
1.2 Reaction Progress Coordinates
The raw $(r,p)$ plane contains an unphysical region where both RMSD distances are simultaneously small, which has no counterpart in configuration space for non-trivial rearrangements. We apply a rigid rotation that decomposes the RMSD plane into reaction progress $s$ and orthogonal deviation $d$, analogous to the path collective variable decomposition [6]. Given the first and last path images with projected coordinates $(r_{0},p_{0})$ and $(r_{N},p_{N})$, we define the unit tangent $(\hat{s}_{r},\hat{s}_{p})$ along the path direction and the normal $(\hat{d}_{r},\hat{d}_{p})=(-\hat{s}_{p},\hat{s}_{r})$. The rotated coordinates are then:
$s_{i}$ $=(r_{i}-r_{0})\,\hat{s}_{r}+(p_{i}-p_{0})\,\hat{s}_{p}$ (4)
$d_{i}$ $=(r_{i}-r_{0})\,\hat{d}_{r}+(p_{i}-p_{0})\,\hat{d}_{p}$ (5)
where $s$ measures progress along the reaction path and $d$ measures perpendicular deviation from it, both in Angstroms. The free energy literature [6] defines path progress and orthogonal distance via an exponential softmin over $N$ reference frames, $s=\sum_{i}i\,e^{-\lambda R_{i}}/\sum_{i}e^{-\lambda R_{i}}$ and $z=-(1/\lambda)\ln\sum_{i}e^{-\lambda R_{i}}$, with a smoothing parameter $\lambda$. That formulation targets enhanced sampling methods (metadynamics, umbrella sampling) where smooth, differentiable collective variables are required for biasing forces at finite temperature. For post-hoc visualization of a zero-temperature discrete path, the linear rotation suffices as it is parameter-free, exact at the endpoints, and directly invertible. Both decompositions separate progress along a path from deviation orthogonal to it. All figures in this work use this $(s,d)$ frame.
Unlike principal component analysis (PCA) [18] or t-distributed stochastic neighbor embedding (t-SNE) [33] applied to covariance matrices [28, 9], which require a priori selection of descriptors (e.g., bond lengths, angles, SOAP vectors [8]), RMSD-based projection operates directly on Cartesian coordinates without feature engineering. PCA-based methods popular in machine learning interatomic potential (MLIP) communities presuppose that the dominant variance directions align with chemically meaningful coordinates, an assumption that frequently fails for complex rearrangements involving concerted bond breaking and formation. Similarly, manifold learning techniques like t-SNE and UMAP [27] optimize for local neighborhood preservation but lack the absolute geometric reference frame necessary for quantitative cross-method comparison. Sketchmap [32] attempts to preserve both local and global distance structure, but requires on the order of $10^{4}-10^{5}$ samples to learn a reliable mapping, and even then does not guarantee metric preservation of absolute distances. A single NEB calculation produces $\approx 10^{3}$ geometries at most. Our endpoint-distance coordinates provide a universal, reaction-agnostic metric from exactly these $10^{3}$ points, with no training phase, no descriptor selection, and no presupposed chemical intuition about the reaction mechanism.
@@ 1.2 Reaction Progress Coordinates
For systems with well-established collective variables (Ramachandran angles for peptides, donor-acceptor distances for proton transfer), such coordinates provide more direct physical insight [23]. However, the vast majority of transition state searches lack such convenient coordinates. Surface reactions, where an adatom hops into a subsurface site on a metallic slab, or catalytic processes involving concerted rearrangements of adsorbates, have no obvious low-dimensional collective variable. Constructing a two-dimensional free energy surface for such systems requires extensive sampling and a priori knowledge of the relevant degrees of freedom. In contrast, the RMSD-based projection requires only the Cartesian coordinates already produced by the calculation. Unlike free energy surface methods that mandate pre-specification of reaction coordinates, our approach visualizes arbitrary path based methods post hoc, requiring only the endpoints. This eliminates the circular dependency where one must already understand the mechanism to choose coordinates capable of revealing it.
@@
Our method targets the complementary regime to global PES exploration techniques. It provides geometric validation for individual optimization trajectories ($\approx 10^{3}$ samples per NEB calculation) rather than exhaustive basin catalogs requiring $\approx 10^{6}-10^{9}$ samples. Most single NEB calculations involve only $\approx 10^{3}$ geometries concentrated along a one-dimensional path connecting a single reactant-product pair. This sampling is insufficient and inappropriate for techniques designed to map complete basin connectivity or construct disconnectivity graphs [34]. The framework instead asks whether this specific optimization found a physically reasonable barrier topology. For MLIP validation, where the goal is verifying that a machine-learned potential reproduces the correct transition pathway rather than exhaustively cataloging all possible pathways, this geometric diagnostic fills a gap that scalar comparisons cannot.
@@ 1.3 Energy Landscape Projection
Configuration space Standard: 1D profile This work: 2D projectionPES: $E(\mathbf{R})$$\mathbf{R}\in\mathbb{R}^{3N-1}$Many configurationscan share the sameprojected coordinatesss = cumulative distance(path-dependent scalar)E(s)E(s)Loses: geometry, basin topology,cross-method comparison(r,p)(r,p) = RMSD to endpoints(intrinsic, path-independent)Rotate $(r,p)\to(s,d)$progress + deviationE(s,d)E(s,d)Loses: 3N33N{-}3 dimensionsRetains: endpoint distances,relative geometryproject ontocumulative path length(3N113N{-$\to 1$)project ontoRMSD distances to $R$, PP($3N{-}1\to 2$)
Figure 2: Dimensional reduction from the full 3N13N{-}1 configuration space to 1D (standard reaction coordinate $s$, left) versus 2D RMSD projection (this work, right). Both maps are many-to-one, and distinct Cartesian configurations can share the same projected coordinates. The 1D projection loses all geometric context. The 2D projection retains endpoint distances for cross-method comparison and landscape visualization, at the cost of 3N33N{-}3 unconstrained degrees of freedom.
The discrete optimization steps provide a sparse, unstructured sampling of the 2D $(s,d)$ domain (Figure 2). To visualize slices of the underlying potential energy landscape $E(s,d)$, in a meaningful way to compare data from disparate calculations, we construct a continuous projection approximating the potential energy surface. Both the 1D and 2D projections are lossy, since multiple distinct configurations in $\mathbb{R}^{3N}$ can map to the same projected coordinate. The 2D projection retains strictly more information than the 1D reaction coordinate, but it is not and cannot be the full PES.
@@
Only the tangential force component $F_{\parallel}$ enters this construction. The full Cartesian force $\mathbf{F}_{i}\in\mathbb{R}^{3N}$ contains orthogonal components that encode local curvature of the true PES. Projecting these into the $(r,p)$ subspace would require the Jacobian $\partial d_{\text{RMSD}}/\partial\mathbf{R}$, which is not analytically available. The RMSD involves a joint minimization over rotations $\mathbf{Q}\in SO(3)$ and permutations $\mathbf{\Pi}$, and the resulting map from $\mathbb{R}^{3N}$ to $\mathbb{R}^{2}$ is piecewise smooth at best, with non-differentiable boundaries wherever the optimal permutation assignment changes. This is not a software limitation but a mathematical property of the permutation-invariant RMSD itself. The tangential projection is therefore the consistent choice without additional approximations.
The omission of orthogonal force components means the interpolated surface cannot reproduce the curvature perpendicular to the path. In the full \$3N\$-dimensional PES, MEP images are by definition minima in the hyperplane orthogonal to the path; the 2D projection has no mechanism to enforce or verify this property. The tangential projection nonetheless carries non-trivial 2D information because the tangent direction $(\tau_{r},\tau_{p})$ varies along the path. Force information is distributed across both projected coordinates in a way that scalar reaction coordinate plots cannot access.
The interpolated surface is a projected slice of the full \$3N\$-dimensional PES. Many distinct Cartesian configurations map to the same $(r,p)$ point. The interpolation represents the conditional expectation of the energy given the RMSD coordinates, marginalizing over the unconstrained 3N33N-3 degrees of freedom. Features of the projected surface, such as apparent alternative pathways in the color map, do not necessarily correspond to physical pathways in the full configuration space. Similarly, the property that MEP images are minima in the orthogonal hyperplane holds in the full-dimensional PES but has no analogue in a two-dimensional lossy projection. The interpolation is reliable near the sampled data and degrades with distance, as with any regression model.
@@
The hyperparameter optimization follows the subset optimization [13] concept from the OT-GPD; here the points from the final path are used to calculate the length and noise scales, which are subsequently applied while fitting the entire surface. This subsampled approach reduces computational cost from $O(N^{3})$ to $O(n^{3})$ where $n\ll N$ (typically $n\approx 20$ vs $N\approx 500-2000$), while simultaneously improving robustness by isolating hyperparameter learning from transient optimization dynamics.
The GP posterior naturally provides a pointwise variance estimate. We overlay variance contours on the energy surface to delineate data-supported regions from extrapolated ones. Near the sampled path, the posterior variance is small; it grows with distance from the data and serves as a built-in reliability indicator for the interpolation.
For systems producing more than approximately $10^{3}$ gradient-augmented observations (e.g., periodic slab calculations with many NEB steps), the $O(N^{3})$ cost of the full GP becomes prohibitive. We use a Nystrom low-rank approximation [35] adapted to the gradient-enhanced setting. Given $M\ll N$ inducing points $\{\mathbf{x}_{m}\}_{m=1}^{M}$ selected from the training set, the augmented kernel matrix $\mathbf{K}_{NN}$ (of size $N(D+1)\times N(D+1)$, where $D=2$ is the input dimension) is approximated as
$\tilde{\mathbf{K}}_{NN}\approx\mathbf{K}_{NM}\,\mathbf{K}_{MM}^{-1}\,\mathbf{K}_{MN}$ (12)
where $\mathbf{K}_{NM}$ and $\mathbf{K}_{MM}$ are the cross and inducing kernel matrices, each containing the value, gradient, and Hessian blocks of the IMQ kernel. The predictive mean becomes $\hat{f}(\mathbf{x}_{*})=\mathbf{k}_{*M}\,\boldsymbol{\alpha}_{M}$ with weights obtained from
$\boldsymbol{\alpha}_{M}=\mathbf{L}^{-\top}\left(\mathbf{S}^{-1}\,\mathbf{V}\,\mathbf{y}_{\text{full}}\right),\quad\mathbf{S}=\mathbf{V}\mathbf{V}^{\top}+\sigma^{2}\mathbf{I}_{M(D+1)}$ (13)
where $\mathbf{L}$ is the Cholesky factor of $\mathbf{K}_{MM}+\sigma^{2}\mathbf{I}$ and $\mathbf{V}=\mathbf{L}^{-1}\mathbf{K}_{MN}$. Inducing points are selected from the last converged NEB path(s), which concentrate observations near the final MEP geometry. This reduces the dominant cost from $O(N^{3})$ to $O(NM^{2})$ (typically $M=300$, $N>5000$) and makes the method applicable to crystalline systems.
The resulting interpolated surface enables direct overlay of reference structures (e.g., DFT-optimized saddle points) onto MLIP-generated landscapes. One can then assess whether a potential captures qualitatively correct barrier topology even when geometric displacements occur. This is impossible with 1D energy profiles, where the reaction coordinate axis itself depends on path geometry.